The board of a large corporation is composed of 7 females and 5 males. The board need to obtain a two-thirds majority to update the organization’s strategic alignment. Statistics have shown that females voted in favor of the proposal for updating the strategic alignment 70% of the time, and males 60% of the time. Develop a probability tree to determine the probability that the strategic alignment of the corporation will be updated.

From the analysis, will the plan be passed?
Answer #1

In favour (5/12*0.6 =0.25) Males (5/12) Against (5/12*0.4 = 0.167) 0 In favour (7/12*0.7= 0.408) Females (7/12) Against (7/12

The probability of updating the strategic alignment = 0.25+0.408= 0.658

                                                                                                     65.8% chance

If you liked the answer, please give me an upvote, it will rally encourage me to write more such answer. Thank you.